A particle travels along a straight line with a constant acceleration. When, s = 4 ft, v = 3 ft/s and when

s = 10 ft, v = 8 ft/s. Determine the velocity as a function of position.




1
Expert's answer
2021-09-10T12:17:28-0400

The traveled distance is

"S=S_2-S_1 \\

S=10-4 \\

S=6 \;ft"

Calculate the acceleration using the following formula:

"v^2_1 -v^2_2 = 2as \\

8^2 -3^2 = 2 \times a \times 6 \\

a=\frac{55}{12} \\

a= 4.58 \;ft/s^2"

Calculate the velocity using the following formula:

"a=\frac{dv}{dt} \\

a=(\frac{dv}{ds})(\frac{ds}{dt}) \\

a= v \times (\frac{dv}{ds}) \\

v \times dv = 4.58 \times ds"

Apply the integration on both sides.

"\int vdv = 4.58 \int ds \\

\frac{v^2}{2}=4.58 \times s +C \\

\frac{3^2}{2}= 4.58 \times 4 + C \\

C= -13.82 \\

\frac{v^2}{2}=4.58 \times s -13.82 \\

v^2 = 9.16s -27.6 \\

v = \sqrt{(9.16s-27.6)}"